OpenStudy (anonymous):

A grocer orders apples and oranges at a total cost of $8.39. If apples cost him $0.25 and oranges cost him $0.18 each, how many of each type of fruit did he order?

OpenStudy (anonymous):

you are trying to solve \[25x+18y=839\] there are infinitely many solutions, but your solution needs to be positive integers

OpenStudy (anonymous):

can you give me a solution for this problem?

OpenStudy (anonymous):

sure try 17 apples and 23 oranges and see if it works

OpenStudy (anonymous):

\[17\times 25+23\times 18=839\] i believe

OpenStudy (anonymous):

I have one more problem. A piggy bank contains 24 coins, all of which are nickels, dimes, or quarters. If the total value of the coins is two dollars, what combinations of coins are possible?

OpenStudy (anonymous):

and thank you that works.

OpenStudy (anonymous):

no idea, but i guess you can guess and check for this one where did you get such an annoying problem?

OpenStudy (anonymous):

my honors algebra 2 teacher lol.

OpenStudy (anonymous):

well, i suppose you are supposed to solve for \(y\) and \(z\) in terms of \(x\) then find a value of \(x\) that makes all three positive

OpenStudy (anonymous):

there may even be more than one solution because i think the equations are \(y=-4x+16\) and \(z=3x+8\)

OpenStudy (anonymous):

with \(x,y,z\) as quarters, dimes and nickels respectively lets try it with \(x=1\) then \(y=12\) and \(z=11\) one quarter, 12 dimes, 11 nickels may work

OpenStudy (anonymous):

\[25+120+55=200\] how about that?